OpenStudy (anonymous):

URGENT:FUNDAMENTAL THEOREM OF SIMILARITY!!! A regular hexagon with a side length of 4 cm has an exact area of 24 square root 3 cm squared. What is the exact area of similar hexagon with a side length of 14 cm?

OpenStudy (cwrw238):

the 2 areas are in the ratio 4^2 / 14^2

OpenStudy (anonymous):

can you explain?

OpenStudy (cwrw238):

if ratio of corresponding sides are a / b then areas are in the ratio a^2 / b^2 so in the case of the hexagons, since the sides are in ratio 4 / 14: 4^2 / 14^2 = (24 sqrt3) ^2 / A where A is area of larger hexagon 16 / 196 = 576*3 / A - solve this for A and you have it.